OpenStudy (anonymous):

How do I solve 0.25r-0.125+0.5r=0.5+r

OpenStudy (whpalmer4):

I might start by multiplying everything by 8 to get rid of the pesky decimals. \[8*(0.25r-0.125+0.5r) = 8*(0.5+r)\]\[2r-1+4r = 4+8r\]Now shuffle things around to get all of the r terms on the left, and the numbers on the right\[2r-1+4r-8r+1=4+8r-8r+1\](adding 1 to both sides, subtracting 8r from both sides)\[-2r=5\]\[r=-2.5\] Plug the answer back into the original formula to check it: \[0.25(-2.5)-0.125+0.5(-2.5) = 0.5+(-2.5)\]\[-0.625-0.125-1.25=-2\]\[-2=-2\]Looks like our answer is correct!